OpenStudy (anonymous):

PLEASE help? X and Y intercepts? Brett is has money to spend for a party. He has $22. He wants to buy nachos for $2.75 and drinks for $2.00 each. Use the equation 2x+2.75y=22 What do the x intercept represent? What does the y intercept represent?

OpenStudy (anonymous):

To find the x intercept, sub in 0 for y 2x + 2.75y = 22 2x + 2.75(0) = 22 2x = 22 x = 11 --- the x intercept tells you Brett bought 11 drinks. To find yhe y intercept, sub in 0 for x 2x + 2.75y = 22 2(0) + 2.75y = 22 2.75y = 22 y = 22/2.75 y = 8 --- the y intercept tells you Brett bought 8 orders of nachos

OpenStudy (anonymous):

@kelliegirl33 so what does the x and y intercept represent in the equation ?

OpenStudy (anonymous):

x intercept represents the number of drinks he bought and the y intercept represents the number of nachos he bought